# Beyond Synthetic Benchmarks: Evaluating LLM Performance on Real-World Class-Level Code Generation

### v1

@@ 9.1.5. Prompt Engineering and Generation Settings
We used a standardized prompt across all models with consistent generation settings to ensure fair comparison, but this does not reflect optimal prompting strategies for each model. Some models may benefit from model-specific prompt formats, few-shot examples (yuenprompting, ), or interactive refinement strategies. However, the finding that all seven models show similar error patterns suggests that fundamental capabilities rather than prompt optimization drive performance. The goal was to assess base capabilities under standardized conditions rather than optimizing each model individually, reflecting realistic deployment scenarios.
@@ 10.3. Construct Validity
Construct validity concerns whether measurements accurately capture intended constructs. We use pass rate as the primary correctness metric, which has limitations: it depends entirely on test suite quality, measures behavioural rather than structural correctness, and doesn’t distinguish between semantically equivalent implementations. However, pass rate is standard in code generation research (yeo2024framework, ; hendrycks2021measuring, ; yuenprompting, ), provides an objective functional correctness measurement through execution, and is validated across multiple test cases per class. Our qualitative analysis examined test suites and failure modes, confirming that pass rates meaningfully distinguish correct implementations from failures.
@@ 11. Conclusion
Our studies reveal that documentation and retrieval provide complementary but modest improvements. This implies that the deployment of RAG should be done adaptively based on documentation completeness. The concentrated error patterns point to critical research directions: developing training objectives targeting attribute resolution and type consistency, creating context-aware retrieval systems that filter incompatible dependencies, and building large-scale real-world benchmarks spanning multiple languages and domains. As LLM-based development tools become increasingly common in software engineering, rigorous evaluation on real-world tasks becomes essential. This work establishes that synthetic benchmarks provide false confidence, and the path forward requires both better models with enhanced object-oriented understanding and better evaluation practices through diverse real-world benchmarks. We make our data and scripts available to facilitate replication and future research (githubGitHubMrsumitbdRealClassEvalReplication, ).

### v2

@@ 9.1.5. Prompt Engineering and Generation Settings
We used a standardized prompt across all models with consistent generation settings to ensure fair comparison, but this does not reflect optimal prompting strategies for each model. Some models may benefit from model-specific prompt formats, few-shot examples (yuenprompting), or interactive refinement strategies. However, the finding that all seven models show similar error patterns suggests that fundamental capabilities rather than prompt optimization drive performance. The goal was to assess base capabilities under standardized conditions rather than optimizing each model individually, reflecting realistic deployment scenarios.
@@ 10.3. Construct Validity
Construct validity concerns whether measurements accurately capture intended constructs. We use pass rate as the primary correctness metric, which has limitations: it depends entirely on test suite quality, measures behavioural rather than structural correctness, and doesn’t distinguish between semantically equivalent implementations. However, pass rate is standard in code generation research (yeo2024framework; hendrycks2021measuring; yuenprompting), provides an objective functional correctness measurement through execution, and is validated across multiple test cases per class. Our qualitative analysis examined test suites and failure modes, confirming that pass rates meaningfully distinguish correct implementations from failures.
@@ 11. Conclusion
Our studies reveal that documentation and retrieval provide complementary but modest improvements. This implies that the deployment of RAG should be done adaptively based on documentation completeness. The concentrated error patterns point to critical research directions: developing training objectives targeting attribute resolution and type consistency, creating context-aware retrieval systems that filter incompatible dependencies, and building large-scale real-world benchmarks spanning multiple languages and domains. As LLM-based development tools become increasingly common in software engineering, rigorous evaluation on real-world tasks becomes essential. This work establishes that synthetic benchmarks provide false confidence, and the path forward requires both better models with enhanced object-oriented understanding and better evaluation practices through diverse real-world benchmarks. We make our data and scripts available to facilitate replication and future research (githubGitHubMrsumitbdRealClassEvalReplication).